Sofia and Luis want to prepare an emergency fund. Together, they earn $50,000 per year. They save 10% of their income, and they pay roughly $9,000 in taxes per year. How much should they at least have in their emergency fund?

Please explain how to get the answer!!!!
Answer #1

Annual Income = $50000

Savings = 10% of income = 10% * 50000 = 5000

Taxes Paid = $9000

Living Expenses = Annual Income - Taxes Paid - Savings

= 50000 - 9000 - 5000

=36000

Monthly Living expense = 36000 / 12 = $3000

According to a rule of thumb an emergency fund should at least have 3 months worth of living expenses

Hence, minimum amount in the emergency fund = 3*$3000 = $9000

A conservative view estimates the emergency fund to have a minimum of 6 months worth of living expenses.

In this case minumum amount in the emergency fund = 6*$3000 = $18000